Q3. An urn contains 7 balls, 4 of which are black and the rest white.
(a) Two are drawn together and one is found to be black. What is the probability that the other is also?
[5 marks]
(b) Starting with all 7 balls in the urn, balls are drawn randomly without replacement until all 3 white balls are found. What is the probability that the third white ball is found on the
(i) fourth attempt? (ii) seventh attempt?
[10 marks]
1
Expert's answer
2011-11-22T12:20:36-0500
(a) & there are 3 black balls left
And also 3 white
So p(ball=black)=3/6=0.5
(b) (i) & w-we pick white ball
b- black/
white ball is found on the fourth attempt& so there can be combinations
wwbw p=3/7*2/6*4/5*1/4=1/35
wbww& p=3/7*4/6*2/5*1/4=1/35
bwww p=4/7*3/6*2/5*1/4=1/35
P=3/35
(ii)
The prob. Is absolutely the same as if we pick the last ball,and it should be white
P(ball=white)=3/7